# Understanding the Training and Generalization of Pretrained Transformer for Sequential Decision Making

### v1

@@ Appendix D Proofs
In addition, we note that in our pretrained transformer setting, the joint distribution can be factorized in
@@
That is, the parameter $\theta$ only induces a covariate shift by affecting the marginal distribution of $X$, but the conditional distribution $\mathcal{P}_{Y|X}$ remains the same for all the $\theta$. This exactly matches our setting of pretrained transformer; to see this, different decision functions, $f$ or $\texttt{TF}_{\theta_{m}}$, only differ in terms of the generation of the actions $a_{\tau}$’s in $H_{t}$ (for $\tau=1,...,t-1$) which is the features $X$ in this formulation but the optimal action $a_{t}^{*}$ will only be affected by $X_{t}$ in $H_{t}$.
@@
Back to the context of the pretrained transformer, such a nice property is contingent on two factors: (i) the transformer function class is rich enough to cover $f^{*}$; (ii) there are infinitely many samples/we can use the expected loss. Also, the above argument is connected to the argument of Proposition 4.1.
@@
$\kappa L_{f}(\texttt{TF}_{\hat{\theta}})+(1-\kappa)L_{\texttt{TF}_{\tilde{\theta}}}(\texttt{TF}_{\hat{\theta}})\leq\frac{1}{nT}\mathop{\sum}\limits_{i=1}^{n}\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\hat{\theta}}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)+\sqrt{\frac{\text{Complex}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{4}{h}\right)},$
where $\text{Complex}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})$ is some inherent constant describing the complexity of $\{\texttt{TF}_{\theta}:\theta\in\Theta\}$. Note that when samples follow the generation process (actions are generated by $f$)
@@
$\leq$ $\mathbb{E}_{(H_{t},a_{t}^{*})\sim\mathcal{P}_{\gamma,f}}\left[\mathop{\sup}\limits_{\theta\in\Theta}\Bigg{|}\frac{1}{nT}\mathop{\sum}\limits_{i=1}^{n}\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)-\mathbb{E}_{(H_{t},a_{t}^{*})\sim\mathcal{P}_{\gamma,f}}\left[\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}\right),a_{t}^{*}\right)\right]\Bigg{|}\right]+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{2}{h}\right)}$
@@
$\leq$ $\sqrt{\frac{\text{Complex}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{2}{h}\right)},$ (10)
@@
$\mathop{\sup}\limits_{\theta\in\Theta}\Bigg{|}\frac{1}{nT}\mathop{\sum}\limits_{i=1}^{n}\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)-L_{\texttt{TF}_{\tilde{\theta}}}(\texttt{TF}_{\theta})\Bigg{|}\leq\sqrt{\frac{\text{Complex}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{2}{h}\right)}.$ (11)
@@
$\kappa L_{f}(\texttt{TF}_{\hat{\theta}})+(1-\kappa)L_{\texttt{TF}_{\tilde{\theta}}}(\texttt{TF}_{\hat{\theta}})\leq\frac{1}{nT}\mathop{\sum}\limits_{i=1}^{n}\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\hat{\theta}}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)+\sqrt{\frac{\text{Complex}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{4}{h}\right)}.$
@@
$L(\texttt{TF}_{\hat{\theta}})\leq L_{f}(\texttt{TF}_{\hat{\theta}})+TD\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,f},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right],\ L(\texttt{TF}_{\hat{\theta}})\leq L_{f}(\texttt{TF}_{\hat{\theta}})+TD\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,\texttt{TF}_{\tilde{\theta}}},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right],$
@@
$L(\texttt{TF}_{\hat{\theta}})\leq$ $\hat{L}(\texttt{TF}_{\hat{\theta}})+\sqrt{\frac{\text{Complex}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}$
$+$ $\kappa TD\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,f},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right]+(1-\kappa)TD\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,\texttt{TF}_{\tilde{\theta}}},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right]+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{4}{h}\right)},$
@@
D.4 Examples for the optimal decision function $\texttt{Alg}^{*}$ in Section 4
Here we provide some additional examples in each problem context to generate more intuitions of the optimal decision function $\texttt{Alg}^{*}$ in Section 4.
(1). Multi-armed bandits under the cross-entropy loss: For convenience, we rewrite the best arm that corresponds to environment $\gamma$, i.e., $a^{*}(H_{t},\gamma)=\mathop{\operatorname*{arg\,max}}\limits_{i}\ \mathbb{E}\left[o_{t}|\gamma_{i}\right]$ as $i_{\gamma}$. Then,
@@ D.4 Examples for the optimal decision function $\texttt{Alg}^{*}$ in Section 4
$\texttt{Alg}^{*}(H_{t})$ $=\operatorname*{arg\,min}_{\sum_{i=1}^{k}p_{i}^{t}=1}\int_{\gamma}-\log\left(p_{i_{\gamma}}^{t}\right)\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P}_{\gamma}=\operatorname*{arg\,min}_{\sum_{i=1}^{k}p_{i}^{t}=1}\ \sum_{j=1}^{k}\left(-\log\left(p_{j}^{t}\right)\right)\int_{\left\{\gamma:\ i_{\gamma}=j\right\}}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P}_{\gamma}.$
@@
By nothing $\int_{\left\{\gamma:\ i_{\gamma}=j\right\}}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P}_{\gamma}$ is the probability such that action $j$ is the optimal action conditional on $H_{t}$ (i.e., the posterior distribution of the optimal action), we are indeed minimizing the cross-entropy of the posterior distribution of the optimal action relative to the decision variables $p_{j}$. Thus, the optimal solution $p^{*}_{j}=\int_{\left\{\gamma:\ i_{\gamma}=j\right\}}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P}_{\gamma}$ for each $j$ and $\texttt{Alg}^{*}(H_{t})$ behaves as the posterior sampling.
(2). Linear bandits under the squared $L_{2}$ loss: We rewrite the best arm that corresponds to environment $\gamma$, i.e., $a^{*}(H_{t},\gamma)=\mathop{\operatorname*{arg\,max}}\limits_{a\in\mathcal{A}}\ w_{\gamma}^{T}a$ as $a_{(\gamma)}$. Then
$\texttt{Alg}^{*}(H_{t})$ $=\operatorname*{arg\,min}_{a\in\mathcal{A}}\int_{\gamma}||a-a^{*}(H_{t},\gamma)||^{2}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}=\operatorname*{arg\,min}_{a\in\mathcal{A}}\int_{\gamma}||a-a_{(\gamma)}||^{2}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}$
$=\frac{\int_{\gamma}a_{(\gamma)}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}}{\int_{\gamma}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}}=\int_{\gamma}a_{(\gamma)}\ \mathrm{d}\mathcal{P}\left(\gamma|H_{t}\right),$
where $\mathcal{P}\left(\gamma|H_{t}\right)$ is the posterior distribution of $\gamma$. Notice that the second last equation holds by taking the derivative of $a$ and applying the first-order condition. Hence $\texttt{Alg}^{*}(H)$ is the posterior averaging.
(3). Newsvendor problem under the $L_{1}$ loss. Let
$a^{*}(H_{t},\gamma)=\mathop{\operatorname*{arg\,max}}\limits_{a\in\mathbb{R}}\ r(X_{t},a)=\mathop{\operatorname*{arg\,max}}\limits_{a\in\mathbb{R}}\ \mathbb{E}\left[-h\cdot(a_{t}-o_{t})^{+}-l\cdot(o_{t}-a_{t})^{+}|X_{t},a_{t}\right]\triangleq a_{(\gamma)}.$
@@
Then by zero-subgradient condition, we can conclude that $\texttt{Alg}^{*}(H_{t})=\mathop{\operatorname*{arg\,min}}\limits_{a\in\mathbb{R}}\int_{\gamma}|a-a_{(\gamma)}|\ \mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}$ satisfies $\int_{\left\{\gamma:\ a\leq a_{(\gamma)}\right\}}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}=\int_{\left\{\gamma:\ a\geq a_{(\gamma)}\right\}}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}$. Divide both sides of this equation by $\int_{\gamma}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}$, we can conclude that
$\int_{\left\{\gamma:\ a\leq a_{(\gamma)}\right\}}\mathrm{d}\mathcal{P}\left({\gamma}|H_{t}\right)=\int_{\left\{\gamma:\ a\geq a_{(\gamma)}\right\}}\mathrm{d}\mathcal{P}\left({\gamma}|H_{t}\right),$
@@
(1). For the linear bandits problem, we consider the following toy model. Consider two environments $\left\{\gamma_{1},\gamma_{2}\right\}$, the prior distribution $\mathcal{P}_{\gamma}$ is given by $\mathcal{P}(\gamma_{1})=\mathcal{P}(\gamma_{2})=\frac{1}{2}$. There is no context, the action space is 2-dimensional. Condition on $\gamma,a_{1},o_{1},\ldots,a_{t}$, the conditional distribution of $o_{t}$ (which contains the reward $R_{t}$ only) is the normal distribution $N\left(w_{\gamma}^{T}a_{t},1\right)$, where $\omega_{\gamma}$ will be given later.
Given $\gamma$ and $H_{t}$, $a_{t}$ is obtained by the posterior averaging
$a_{t}=\texttt{Alg}^{*}(H_{t})=\operatorname*{arg\,min}_{a\in\mathcal{A}}\int_{\gamma}||a-a^{*}(H_{t},\gamma)||^{2}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}.$
Denote $a^{*}(H_{t},\gamma)=\mathop{\operatorname*{arg\,max}}\limits_{a\in\mathcal{A}}\ w_{\gamma}^{T}a$ by $a_{(\gamma)}$. Let $w_{\gamma_{1}}=(1,0)^{T}$, $w_{\gamma_{2}}=(0,1)^{T}$, and $\mathcal{A}=[-1,1]\times[-1,1]$, $a_{(\gamma_{1})}=(1,0)^{T}$, $a_{(\gamma_{2})}=(0,1)^{T}$.
Observe that by part (b) in Example 4.2 (see also D.4 for more details), we have
$a_{t}=\texttt{Alg}^{*}(H_{t})=\mathcal{P}\left(\gamma_{1}|H_{t}\right)\cdot a_{(\gamma_{1})}+\mathcal{P}\left(\gamma_{2}|H_{t}\right)\cdot a_{(\gamma_{2})},$
@@
where $\mathcal{P}\left(\gamma|H_{t}\right)$ is the posterior of $\gamma$ condition on $H_{t}$, and is given by
@@
$\mathcal{P}\left(\gamma_{i}|H_{t}\right)=\frac{\prod_{\tau=1}^{t-1}\mathbb{P}_{i}(o_{\tau}|a_{\tau})}{\sum_{i=1}^{2}\prod_{\tau=1}^{t-1}\mathbb{P}_{i}(o_{\tau}|a_{\tau})},\quad i=1,2.$
@@
We now use induction to show $\mathcal{P}\left(\gamma_{i}|H_{t}\right)=\frac{1}{2}$ for any $t$, thereby having $a_{t}=\frac{1}{2}a_{(\gamma_{1})}+\frac{1}{2}a_{(\gamma_{2})}=\left(\frac{1}{2},\frac{1}{2}\right)^{T}$ does not change with respect to $t$.
@@
Step 1. Since $\mathcal{P}\left(\gamma_{i}|H_{0}\right)=\mathcal{P}\left(\gamma_{i}\right)=\frac{1}{2}$, we have $a_{1}=\frac{1}{2}a_{(\gamma_{1})}+\frac{1}{2}a_{(\gamma_{2})}=(\frac{1}{2},\frac{1}{2})^{\top}$, and the conclusion holds for $t=1$.
@@
Step 2. Now assume the conclusion holds for $t$, i.e., $\mathcal{P}\left(\gamma_{i}|H_{t}\right)=\frac{1}{2}$, and $a_{t}=\left(\frac{1}{2},\frac{1}{2}\right)^{T}$. Since $w_{\gamma_{1}}^{T}a_{t}=w_{\gamma_{2}}^{T}a_{t}=\frac{1}{2}$, we have $\mathbb{P}_{i}(o_{t}|a_{t})=\frac{1}{\sqrt{2\pi}}\exp\left(-\frac{\left(o_{t}-\frac{1}{2}\right)^{2}}{2}\right)$. Observe that
$\mathcal{P}(\gamma_{i}|H_{t+1})=\frac{\mathcal{P}(\gamma_{i},H_{t+1})}{\sum_{i^{\prime}=1}^{2}\mathcal{P}(\gamma_{i^{\prime}},H_{t+1})}=\frac{\mathcal{P}(H_{t})\mathcal{P}(\gamma_{i}|H_{t})\mathbb{P}_{i}(o_{t}|a_{t})}{\sum_{i^{\prime}=1}^{2}\mathcal{P}(H_{t})\mathcal{P}(\gamma_{i^{\prime}}|H_{t})\mathbb{P}_{i^{\prime}}(o_{t}|a_{t})}=\frac{\mathbb{P}_{i}(o_{t}|a_{t})}{\sum_{i^{\prime}=1}^{2}\mathbb{P}_{i^{\prime}}(o_{t}|a_{t})}=\frac{1}{2},$
which implies $a_{t+1}=\left(\frac{1}{2},\frac{1}{2}\right)^{T}$, and the conclusion holds for $t+1$.
Therefore, by the principle of induction, the conclusion holds. Then the regret for $\gamma$ is
@@
$\text{Regret}(\texttt{Alg}^{*};\gamma)=\mathbb{E}\left[\sum_{t=1}^{T}r(X_{t},a_{t}^{*})-r(X_{t},a_{t})\right]=\frac{1}{2}T,$
which completes the first part of the proof.

### v2

@@ Appendix D Proofs
In addition, we note that in our pre-trained transformer setting, the joint distribution can be factorized in
@@
That is, the parameter $\theta$ only induces a covariate shift by affecting the marginal distribution of $X$, but the conditional distribution $\mathcal{P}_{Y|X}$ remains the same for all the $\theta$. This exactly matches our setting of pre-trained transformer; to see this, different decision functions, $f$ or $\texttt{TF}_{\theta_{m}}$, only differ in terms of the generation of the actions $a_{\tau}$’s in $H_{t}$ (for $\tau=1,...,t-1$) which is the features $X$ in this formulation but the optimal action $a_{t}^{*}$ will only be affected by $X_{t}$ in $H_{t}$.
@@
Back to the context of the pre-trained transformer, such a nice property is contingent on two factors: (i) the transformer function class is rich enough to cover $f^{*}$; (ii) there are infinitely many samples/we can use the expected loss. Also, the above argument is connected to the argument of Proposition 4.1.
@@
$\kappa L_{f}(\texttt{TF}_{\hat{\theta}})+(1-\kappa)L_{\texttt{TF}_{\tilde{\theta}}}(\texttt{TF}_{\hat{\theta}})\leq$ $\frac{1}{nT}\mathop{\sum}\limits_{i=1}^{n}\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\hat{\theta}}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)$
$+\sqrt{\frac{\text{Comp}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{4}{h}\right)},$
where $\text{Comp}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})$ is some inherent constant describing the complexity of $\{\texttt{TF}_{\theta}:\theta\in\Theta\}$. Note that when samples follow the generation process (actions are generated by $f$)
@@
$\leq$ $\mathbb{E}_{(H_{t},a_{t}^{*})\sim\mathcal{P}_{\gamma,f}}\left[\mathop{\sup}\limits_{\theta\in\Theta}\Bigg{|}\frac{1}{nT}\mathop{\sum}\limits_{i=1}^{n}\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)-\mathbb{E}_{(H_{t},a_{t}^{*})\sim\mathcal{P}_{\gamma,f}}\left[\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}\right),a_{t}^{*}\right)\right]\Bigg{|}\right]$ (10)
$+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{2}{h}\right)}$
@@
$\leq$ $\sqrt{\frac{\text{Comp}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{2}{h}\right)},$ (11)
@@
$\mathop{\sup}\limits_{\theta\in\Theta}\Bigg{|}\frac{1}{nT}\mathop{\sum}\limits_{i=1}^{n}\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)-L_{\texttt{TF}_{\tilde{\theta}}}(\texttt{TF}_{\theta})\Bigg{|}\leq\sqrt{\frac{\text{Comp}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{2}{h}\right)}.$ (12)
@@
$\kappa L_{f}(\texttt{TF}_{\hat{\theta}})+(1-\kappa)L_{\texttt{TF}_{\tilde{\theta}}}(\texttt{TF}_{\hat{\theta}})\leq$ $\frac{1}{nT}\mathop{\sum}\limits_{i=1}^{n}\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\hat{\theta}}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)$
$+\sqrt{\frac{\text{Comp}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{4}{h}\right)}.$
@@
$L(\texttt{TF}_{\hat{\theta}})\leq L_{f}(\texttt{TF}_{\hat{\theta}})+TD\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,f},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right],$
$L(\texttt{TF}_{\hat{\theta}})\leq L_{f}(\texttt{TF}_{\hat{\theta}})+TD\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,\texttt{TF}_{\tilde{\theta}}},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right],$
@@
$L(\texttt{TF}_{\hat{\theta}})\leq$ $\hat{L}(\texttt{TF}_{\hat{\theta}})+\sqrt{\frac{\text{Comp}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}$
$+\kappa TD\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,f},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right]$
$+(1-\kappa)TD\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,\texttt{TF}_{\tilde{\theta}}},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right]+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{4}{h}\right)},$
@@
D.4 Proof of Example 4.2
Proof.
Since $\texttt{Alg}^{*}$ is a function of any possible history $H_{t}$, we omit the subscrpit $t$ for the optimal actions for notation simplicity, and use $a^{(\gamma)*}$ as the optimal action of $\gamma$.
  • Posterior sampling under the cross-entropy loss. We assume that $\mathcal{A}=\{a_{1},\ldots,a_{J}\}$ and the output of $\texttt{Alg}^{*}$ is a probability vector $(p_{1},\ldots,p_{J})$ such that $\sum_{j=1}^{J}p_{j}=1$, meaning the probability of choosing each action. Then
@@
    $\texttt{Alg}^{*}(H)$ $=\operatorname*{arg\,min}_{\sum_{j=1}^{J}p_{j}=1}\int_{\gamma}-\log\left(p_{a^{(\gamma)*}}\right)\mathbb{P}(H|\gamma)\mathrm{d}\mathcal{P}_{\gamma}$
    $=\operatorname*{arg\,min}_{\sum_{j=1}^{J}p_{j}=1}\ \sum_{j=1}^{J}\left(-\log\left(p_{j}\right)\right)\int_{\left\{\gamma:\ a^{(\gamma)*}=j\right\}}\mathbb{P}(H|\gamma)\mathrm{d}\mathcal{P}_{\gamma}.$
@@
    By nothing $\int_{\left\{\gamma:\ a^{(\gamma)*}=j\right\}}\mathbb{P}(H|\gamma)\mathrm{d}\mathcal{P}_{\gamma}$ is the probability such that action $j$ is the optimal action conditional on $H$ (i.e., the posterior distribution of the optimal action), we are indeed minimizing the cross-entropy of the posterior distribution of the optimal action relative to the decision variables $p_{j}$. Thus, the optimal solution $p^{*}_{j}=\int_{\left\{\gamma:\ a^{(\gamma)*}=j\right\}}\mathbb{P}(H|\gamma)\mathrm{d}\mathcal{P}_{\gamma}$ for each $j$ and $\texttt{Alg}^{*}(H_{t})$ behaves as the posterior sampling.
  • Posterior averaging under the squared loss. By definition,
    $\texttt{Alg}^{*}(H)$ $=\operatorname*{arg\,min}_{a\in\mathcal{A}}\int_{\gamma}||a-a^{(\gamma)*}||^{2}_{2}\mathbb{P}(H|\gamma)\mathrm{d}\mathcal{P_{\gamma}}$
    $=\frac{\int_{\gamma}a^{(\gamma)*}\mathbb{P}(H|\gamma)\mathrm{d}\mathcal{P_{\gamma}}}{\int_{\gamma}\mathbb{P}(H|\gamma)\mathrm{d}\mathcal{P_{\gamma}}}$
    $=\mathbb{E}_{\gamma}[a^{(\gamma)*}|H]$
    where the second line is by the first order condition.
  • Posterior median under the absolute loss. By definition,
    $\texttt{Alg}^{*}(H)$ $=\operatorname*{arg\,min}_{a\in\mathcal{A}}\int_{\gamma}|a-a^{(\gamma)*}|\mathbb{P}(H|\gamma)\mathrm{d}\mathcal{P_{\gamma}}.$
    Then by zero-subgradient condition, we can conclude that $\texttt{Alg}^{*}(H)$ satisfies
    $\int_{\left\{\gamma:\ \texttt{Alg}^{*}(H)\leq a^{(\gamma)*}\right\}}\mathbb{P}(H|\gamma)\mathrm{d}\mathcal{P_{\gamma}}=\int_{\left\{\gamma:\ \texttt{Alg}^{*}(H)\geq a^{(\gamma)*}\right\}}\mathbb{P}(H|\gamma)\mathrm{d}\mathcal{P_{\gamma}}.$
@@
    Divide both sides of this equation by $\int_{\gamma}\mathbb{P}(H|\gamma)\mathrm{d}\mathcal{P_{\gamma}}$, we can conclude that
    $\int_{\left\{\gamma:\ \texttt{Alg}^{*}(H)\leq a^{(\gamma)*}\right\}}\mathrm{d}\mathcal{P}\left({\gamma}|H\right)=\int_{\left\{\gamma:\ \texttt{Alg}^{*}(H)\geq a^{(\gamma)*}\right\}}\mathrm{d}\mathcal{P}\left({\gamma}|H\right)$
@@
For the linear bandits problem, we consider the following example. Suppose we have two environments, ${\gamma_{1},\gamma_{2}}$, each with standard normal distributed noise and unknown parameters $w_{\gamma_{1}}=(1,0)$ and $w_{\gamma_{2}}=(0,1)$, respectively. Let the action space be $\mathcal{A}=[-1,1]\times[-1,1]$. The optimal actions for the two environments are $a^{(1)*}=(1,0)$, $a^{(2)*}=(0,1)$. We assume the prior distribution $\mathcal{P}_{\gamma}$ is given by $\mathcal{P}_{\gamma}(\gamma_{1})=\mathcal{P}_{\gamma}(\gamma_{2})=\frac{1}{2}$.
Then if $\texttt{Alg}^{*}$ is the posterior averaging, we have
$a_{t}=\texttt{Alg}^{*}(H_{t})=\mathcal{P}_{\gamma}\left(\gamma_{1}|H_{t}\right)\cdot a^{(1)*}+\mathcal{P}_{\gamma}\left(\gamma_{2}|H_{t}\right)\cdot a^{(2)*},$
@@
where the posterior distribution $\mathcal{P}_{\gamma}\left(\gamma_{i}|H_{t}\right)$ is given by
@@
$\mathcal{P}_{\gamma}\left(\gamma_{i}|H_{t}\right)=\frac{\prod_{\tau=1}^{t-1}\mathbb{P}_{\gamma_{i}}(o_{\tau}|a_{\tau})}{\sum_{i^{\prime}=1}^{2}\prod_{\tau=1}^{t-1}\mathbb{P}_{\gamma_{i^{\prime}}}(o_{\tau}|a_{\tau})},\quad i=1,2.$
@@
We now use induction to show $\mathcal{P}_{\gamma}\left(\gamma_{i}|H_{t}\right)=\frac{1}{2}$ for any $t\geq 1$ and $i=1,2$, where $H_{t}$ can be generated by either $\gamma_{1}$ or $\gamma_{2}$. This results in $a_{t}=\frac{1}{2}a^{(1)*}+\frac{1}{2}a^{(2)*}=\left(\frac{1}{2},\frac{1}{2}\right)$, which does not change with respect to $t$ and will cause regret linear in $T$.
@@
Step 1. Since $\mathcal{P}_{\gamma}\left(\gamma_{i}|H_{0}\right)=\mathcal{P}_{\gamma}\left(\gamma_{i}\right)=\frac{1}{2}$, we have $a_{1}=\frac{1}{2}a^{(1)*}+\frac{1}{2}a^{(2)*}=(\frac{1}{2},\frac{1}{2})$, and the conclusion holds for $t=1$.
@@
Step 2. Now assume the conclusion holds for $t$, i.e., $\mathcal{P}_{\gamma}\left(\gamma_{i}|H_{t}\right)=\frac{1}{2}$, and $a_{t}=\left(\frac{1}{2},\frac{1}{2}\right)$. Since $w_{\gamma_{1}}^{T}a_{t}=w_{\gamma_{2}}^{T}a_{t}=\frac{1}{2}$, we have $\mathbb{P}_{\gamma_{i}}(o_{t}|a_{t})=\frac{1}{\sqrt{2\pi}}\exp\left(-\frac{\left(o_{t}-\frac{1}{2}\right)^{2}}{2}\right)$ for $i=1,2$. Observe that
$\mathcal{P}_{\gamma}(\gamma_{i}|H_{t+1})=\frac{\mathbb{P}_{\gamma_{i}}(o_{t+1}|a_{t+1})\mathcal{P}_{\gamma}(\gamma_{i}|H_{t})}{\sum_{i^{\prime}=1}^{2}\mathbb{P}_{\gamma_{i^{\prime}}}(o_{t+1}|a_{t+1})\mathcal{P}_{\gamma}(\gamma_{i^{\prime}}|H_{t})}=\frac{1}{2},$
which implies $a_{t+1}=\left(\frac{1}{2},\frac{1}{2}\right)$, and the conclusion holds for $t+1$.
Thus, the conclusion holds for all $t\geq 1$. Then the regret is
@@
$\text{Regret}(\texttt{Alg}^{*};\gamma_{i})=\mathbb{E}\left[\sum_{t=1}^{T}r(X_{t},a_{t}^{*})-r(X_{t},a_{t})\right]=\frac{1}{2}T$
for $i=1,2$.